# Nonlocal characteristics of two-qubit gates and their argand diagrams

### v1

@@ Abstract
In this paper, we show the usefulness of the chords present in the argand diagram of squared eigenvalues of nonlocal part of two-qubit gates to study their nonlocal characteristics. We discuss about the criteria for perfect entanglers to transform a pair of orthonormal product states into a pair of orthonormal maximally entangled states. Perfect entanglers with a chord passing through origin can do such a transformation. We show that all the perfect entanglers with at least one chord passing through origin form five different planes of Weyl chamber. Perfect entanglers represented by each plane have unique entangling characteristics. We also provide the conditions for a perfect entangler not present in one of those five planes to transform a pair of orthonormal product states into orthonormal maximally entangled states. Finally, we show that similar to entangling power, gate typicality can also be described using the chords present in the argand diagram. For each chord describing the entangling power there exists a chord describing the gate typicality. We show the geometrical relation between the two sets of chords.
@@ I Introduction
Understanding the geometry and nonlocal characteristics of two-qubit gates is essential as their role in quantum computation is vital [1, 2, 3]. Over two decades, two-qubit gates have been studied vastly. Nonlocal characteristics of two-qubit gates are invariant under local operations. A complete set of local invariants of two-qubit gates were obtained in [4]. The geometry of local equivalence classes of two-qubit gates was studied in detail in [5]. A measure for operator entanglement was introduced in [6]. Nonlocal characteristics of two-qubit gates were considered as resources for doing quantum information processing and their quantification was studied in [7]. Entangling power, a quantity to measure the ability of two-qubit gates to generate entanglement, was defined in [8]; its expression in terms of Cartan co-ordinates [9] and local invariants [10] were derived. Gate typicality as complementary to entangling power was introduced [11] and its properties for two-qubit gates were studied [12].
@@
Despite many studies, still there remains unexplored ways to understand the nonlocal characteristics of two-qubit gates. Recently, the chords present in the argand diagram of squared eigenvalues of nonlocal parts of two-qubit gates were shown to describe the ability of two-qubit gates to generate entangled states [13]. This argand diagram was studied in regard to the condition for perfect entanglers [5, 4], operational discrimination of nonlocal part of a two-qubit gate from the nonlocal part of its adjoint [14], and simulation of perfect entanglers [15]. Perfect entanglers are two-qubit gates that can transform some product state into a maximally entangled state. For perfect entanglers, convex hull of squared eigenvalues of their nonlocal part contains zero [5, 4]. Among perfect entanglers, special perfect entanglers (SPEs) can transform orthonormal product basis into orthonormal maximally entangled basis [9]. Recently, it was shown that perfect entanglers with a chord passing through origin in their argand diagram can transform a pair of orthonormal product states into a pair of orthonormal maximally entangled states [13]. Now, there ariases a question that can a perfect entangler without any chords passing through origin transform a pair of orthonormal product states into a pair of orthonormal maximally entangled states?
@@
In this paper, we provide answer to this question. First, we identify the regions of perfect entanglers with at least one chord passing through origin. Next, we derive three sets of orthogonal conditions, in terms of eigenvalues of generators of perfect entanglers, between a pair of product states obtained from the argand diagram of perfect entanglers without any chord passing through origin. A perfect entangler without any chord passing through origin and transforming a pair of orthonormal product states into maximally entangled states should satisfy one of those conditions. We also provide an example for existence of such perfect entanglers. Finally, we show that in addition to entangling power, the chords present in the argand diagram can also be used to quantify the gate typicality and linear entropy [6] of two-qubit gates.
@@ II Background
A two-qubit gate $U\in\text{SU}(4)$ can be decomposed as follows [5].
@@
where $L_{1},~{}L_{2}\in\text{SU}(2)\otimes\text{SU}(2)$ are local parts of $U$ and $U_{d}(c_{x},c_{y},c_{z})$ is the nonlocal part of $U$. Two-qubit gates having the same nonlocal part but differing only in the local parts form a local equivalence class and each local equivalence class of two-qubit gates are geometrically represented as a point of tetrahedron shown in FIG. 1.
@@
Figure 1: Geometry of local equivalence classes of two-qubit gates. Region colored in blue is the region of perfect entanglers. The line in red color represent special perfect entanglers.
@@
$U_{d}(c_{1},c_{2},c_{3})=\dfrac{1}{2}\sum_{j=1}^{4}e^{ih_{j}/2}|\Psi_{j}\rangle\langle\Psi_{j}|,$ (5)
where $|\Psi_{j}\rangle$’s are unnormalized vectors forming the eigen basis of $U_{d}$. It is also called as magic basis [5] and the expressions of unnormalized vectors are given below.
$|\Psi_{1}\rangle=\begin{bmatrix}1\\ 0\\ 0\\ 1\end{bmatrix},~{}|\Psi_{2}\rangle=\begin{bmatrix}0\\ i\\ i\\ 0\end{bmatrix},|\Psi_{3}\rangle=\begin{bmatrix}0\\ 1\\ -1\\ 0\end{bmatrix},~{}|\Psi_{4}\rangle=\begin{bmatrix}i\\ 0\\ 0\\ -i\end{bmatrix}.$ (6)
@@
$c_{x}+c_{y}\geq\dfrac{\pi}{2},$
$c_{y}\pm c_{z}\leq\dfrac{\pi}{2}.$ (7)
The region of perfect entanglers (colored in blue) is shown in FIG. 1. Special perfect entanglers which are perfect entanglers with maximum entangling power [9] is represenetd by the red color line in FIG. 1. For perfect entanglers, the convex hull of squared eigenvalues of $U_{d}$ encloses zero. The argand diagram of a typical perfect entangler (non perfect entangler) is shown in FIG. 2a (FIG. 2b).
@@
III Perfect entanglers with chords passing through origin
@@
In this section, we identify the regions of perfect entanglers with at least one chord passing through origin. The chord connecting $e^{ih_{j}}$ and $e^{ih_{k}}$ passes through origin only if $|h_{j}-h_{k}|=\pi$. It can be verified that the equations $|h_{j}-h_{j+1}|=\pi$ with $j=1,2,3$ describe the three boundary planes separating perfect entanglers from non perfect entanglers inside the Weyl chamber [Eq. 7]. These three planes are shown in FIG. 3a. Similarly, the equations $|h_{j}-h_{j+2}|=\pi$ with $j=1,2$ describe the planes $c_{x}\pm c_{z}=\pi/2$ [FIG. 3b]. These two planes together form one of the reflecting plane, $c_{x}+\text{sign}(c_{z})c_{z}=\pi/2$, describing the mirror operation [17]. The equation $|h_{1}-h_{4}|=\pi$ which can be rewritten as $c_{x}-c_{y}=\pi/2$ is satisfied only by the point representing CNOT equivalence class. Thus only the gates represented by the five planes shown in FIG. 3a and FIG. 3b have atleast one chord passing through origin and hence they can transform a pair of orthonormal product states into a pair of orthonormal maximally entangled states.
@@
All the gates represented by a specific plane have unique argand diagram with a specific chord passing through origin. However, there are gates present in more than one plane. SPEs are at the intersection of $|h_{j}-h_{j+2}|=\pi~{}(j=1,2)$ planes. Hence, for SPEs, two chords connecting diametrically opposite points $(e^{ih_{j}}~{}\text{and}~{}e^{ih_{j+2}}~{}\text{with}~{}j=1,2)$ pass through origin and they can transform an orthonormal product basis into orthonormal maximally entangled basis [13, 9]. Among SPEs, the point representing iSWAP equivalence class is also present in $|h_{j}-h_{j+1}|=\pi~{}\text{with}~{}j=1,3$ planes. Similarly, the point representing CNOT equivalence class also satify the conditions, $|h_{2}-h_{3}|=\pi$ and $|h_{1}-h_{4}|=\pi$. Hence, these two equivalence classes have argand diagrams with four chords passing through origin [13].
@@
Similar to SPEs, there are local equivalence classes at the intersection of two planes: between the pairs of planes $(|h_{3}-h_{4}|=\pi,|h_{1}-h_{3}|=\pi)$, $(|h_{2}-h_{3}|=\pi,|h_{1}-h_{3}|=\pi)$, $(|h_{1}-h_{2}|=\pi,|h_{2}-h_{4}|=\pi)$, and $(|h_{2}-h_{3}|=\pi,|h_{2}-h_{4}|=\pi)$. However, for these equivalence classes, the two chords connecting diametrically opposite points do not pass through origin but, only one of them pass through origin. Hence these equivalence classes cannot transform a orthonormal product basis into orthonormal maximally entangled basis.
@@
The point representing $\sqrt{\text{SWAP}}$ equivalence class is common to $|h_{3}-h_{4}|=\pi$, $|h_{1}-h_{3}|=\pi$, and $|h_{2}-h_{3}|=\pi$ planes. Hence, in the argand diagram of $\sqrt{\text{SWAP}}$ equivalence class, three chords from the point $e^{ih_{3}}$ pass through origin. It implies that the remaining three points $e^{ih_{1}}$, $e^{ih_{2}}$, and $e^{ih_{4}}$ coincide with each other and the argand diagram has only three chords. Similarly, it can be verified that the argand diagram of $\sqrt{\text{SWAP}}^{\dagger}$ equivalence class has only three chords from the point $e^{ih_{2}}$ passing through origin. Thus for the gates belonging to these two equivalence classes there exist three different pairs of orthonormal product states that can be transformed into pairs of orthonormal maximally entangled states.
@@
Figure 3: Regions of perfect entanglers with at least one chord passing through origin. In the subfigure (a) the planes colored in meganta, green, and yellow are $|h_{1}-h_{2}|=\pi$, $|h_{2}-h_{3}|=\pi$, and $|h_{3}-h_{4}|=\pi$ planes respectively. The violet line (excluding the end points) represents perfect entanglers without any chord passing through origin but capable of transforming a pair of orthonormal product states into a maximally entangled states. In the subfigure (b), the planes colored in red and blue are $|h_{1}-h_{3}|=\pi$ and $|h_{2}-h_{4}|=\pi$ planes respectively.
IV Pair of orthonormal product states to maximally entangled states
@@
In this section, we consider the perfect entanglers existing in other regions. A typical argand diagram of perfect entanglers is shown in FIG. 2a. In FIG. 2a, the origin is contained in two triangles: the triangle with vertices $e^{ih_{j}}$, $e^{ih_{l}}$, and $e^{ih_{m}}$, and the triangle with vertices $e^{ih_{k}}$, $e^{ih_{l}}$, and $e^{ih_{m}}$. Hence, there exist two sets of weights or co-ordinates such that
@@
with $|\phi_{j}|^{2}+|\phi_{l}|^{2}+|\phi_{m}|^{2}=1$ and $|\phi^{\prime}_{k}|^{2}+|\phi^{\prime}_{l}|^{2}+|\phi^{\prime}_{m}|^{2}=1$. From Eqs. 8 and 9, following two product states can be constructed [5].
$|\Phi_{1}\rangle=|\phi_{j}|e^{-i(a\pi+h_{j}/2)}|\Psi_{j}\rangle+|\phi_{l}|e^{-i(b\pi+h_{l}/2)}|\Psi_{l}\rangle+|\phi_{m}|e^{-i(c\pi+h_{m}/2)}|\Psi_{m}\rangle,$ (10)
@@
$|\Phi_{2}\rangle=|\phi^{\prime}_{k}|e^{-i(d\pi+h_{k}/2}|\Psi_{k}\rangle+|\phi^{\prime}_{l}|e^{-i(f\pi+h_{l}/2)}|\Psi_{l}\rangle+|\phi^{\prime}_{m}|e^{-i(g\pi+h_{m}/2)}|\Psi_{m}\rangle,$ (11)
for some integers $a,b,c,d,f,~{}\text{and}~{}g$. Now we consider the following cases.
Case 1: First, we derive the condition for the existence of a pair of orthonormal product states with respect to a triangle enclosing the origin. For example, in FIG. 2a, we consider the triangle with vertices $e^{ih_{j}}$, $e^{ih_{l}}$, and $e^{ih_{m}}$. A product state corresponding to this triangle is given in Eq. 10. Now we construct another product state with respect to the same triangle, for some integers $a^{\prime},b^{\prime},~{}\text{and}~{}c^{\prime}$, as shown below.
$|\Phi^{\prime}_{1}\rangle=|\phi_{j}|e^{-i(a^{\prime}\pi+h_{j}/2)}|\Psi_{j}\rangle+|\phi_{l}|e^{-i(b^{\prime}\pi+h_{l}/2)}|\Psi_{l}\rangle+|\phi_{m}|e^{-i(c^{\prime}\pi+h_{m}/2)}|\Psi_{m}\rangle.$ (12)
@@
Along with normalization, it can be found that $|\phi_{j}|^{2}=1/2$. Thus, to construct a pair of orthonormal product states with respect to the triangle formed by $e^{ih_{j}}$, $e^{ih_{l}}$, and $e^{ih_{m}}$, one of the weights in Eq. 8 should be $1/2$. The condition for $|\phi_{j}|^{2}=0.5,$ can be written in terms of eigenvalues of $H$ as follows.
$\dfrac{2|\cos h_{k}\sin h_{l}-\cos h_{l}\sin h_{k}|}{|\cos h_{j}(\sin h_{k}-\sin h_{l})+\cos h_{k}(\sin h_{l}-\sin h_{j})+\cos h_{l}(\sin h_{j}-\sin h_{k})|}=1.$ (15)
Similar conditions can also be written for other two weights $|\phi_{l}|^{2}$ and $|\phi_{m}|^{2}$.
Case 2: Now, we consider the case where $e^{ih_{j}}$ coincides with $e^{ih_{k}}$ in FIG. 2a. In this case, we have $|\phi^{\prime}_{k}|=|\phi_{j}|$, $|\phi^{\prime}_{l}|=|\phi_{l}|$, and $|\phi^{\prime}_{m}|=|\phi_{m}|$ in Eqs. 9 and 11. Choosing $(f-b)$ as even (odd) and $(g-c)$ as odd (even), the orthogonalization condition becomes
$|\phi_{l}|^{2}=|\phi_{m}|^{2}.$ (16)
$\dfrac{|\cos h_{m}\sin h_{j}-\cos h_{j}\sin h_{m}|}{|\cos h_{j}\sin h_{l}-\cos h_{l}\sin h_{j}|}=1.$ (17)
Case 3: Now, we derive the condition for the two product states shown in Eqs. 10 and 11 to be orthogonal. These two product states correspond to the two triangles containing the origin in FIG. 2a. The condition for orthogonalization between those two product states can be written as follows.
Above equation can be rearranged after taking square on both sides as follows.
$\dfrac{|\phi_{l}|^{2}}{|\phi_{m}|^{2}}=\dfrac{|\phi^{\prime}_{m}|^{2}}{|\phi^{\prime}_{l}|^{2}}.$ (19)
@@
This condition can be expressed in terms of the eigenvalues of $H$ as follows.
$\dfrac{|\cos h_{m}\sin h_{j}-\sin h_{m}\cos h_{j}|}{|\sin h_{l}\cos h_{j}-\cos h_{l}\sin h_{j}|}=\dfrac{|\sin h_{l}\cos h_{k}-\cos h_{l}\sin h_{k}|}{|\cos h_{m}\sin h_{k}-\sin h_{m}\cos h_{k}|}.$ (20)

### v2

@@ Abstract
In this paper, we show the usefulness of the chords present in the argand diagram of squared eigenvalues of nonlocal part of two-qubit gates to study their nonlocal characteristics. We discuss the criteria for perfect entanglers to transform a pair of orthonormal product states into a pair of orthonormal maximally entangled states. Perfect entanglers with a chord passing through origin can do such a transformation. In the Weyl chamber, we identify the regions of perfect entanglers with at least one chord passing through origin. We also provide the conditions for a perfect entangler without any chord passing through origin to transform a pair of orthonormal product states into orthonormal maximally entangled states. Finally, we show that similar to entangling power, gate typicality can also be described using the chords present in the argand diagram. For each chord describing the entangling power, there exists a chord describing the gate typicality. We show the geometrical relation between the two sets of chords.
@@ I Introduction
Understanding the geometry and nonlocal characteristics of two-qubit gates is essential as their role in quantum computation is vital [1, 2, 3]. Over two decades, two-qubit gates have been studied vastly. Nonlocal characteristics of two-qubit gates are invariant under local operations. A complete set of local invariants of two-qubit gates was obtained in [4]. The geometry of local equivalence classes of two-qubit gates was studied in detail in [5]. A measure for operator entanglement was introduced in [6]. Nonlocal characteristics of two-qubit gates were considered as resources for doing quantum information processing and their quantification were studied in [7]. Entangling power, a quantity to measure the ability of two-qubit gates to generate entanglement, was defined in [8]; its expression in terms of Cartan co-ordinates [9] and local invariants [10] were derived. Gate typicality as complementary to entangling power was introduced [11] and its properties for two-qubit gates were studied [12].
@@
Despite many studies, still there remain unexplored ways to understand the nonlocal characteristics of two-qubit gates. Recently, the chords present in the argand diagram of squared eigenvalues of nonlocal parts of two-qubit gates were shown to describe the ability of two-qubit gates to generate entangled states [13]. This argand diagram was studied in regard to the condition for perfect entanglers [5, 4], operational discrimination of the nonlocal part of a two-qubit gate from the nonlocal part of its adjoint [14], and simulation of perfect entanglers [15]. Perfect entanglers are two-qubit gates that can transform some product state into a maximally entangled state. For perfect entanglers, the convex hull of squared eigenvalues of their nonlocal part contains zero [5, 4]. Among perfect entanglers, special perfect entanglers (SPEs) can transform orthonormal product basis into orthonormal maximally entangled basis [9]. Recently, it was shown that perfect entanglers with a chord passing through origin in their argand diagram can transform a pair of orthonormal product states into a pair of orthonormal maximally entangled states [13]. Now, there arises a question that can a perfect entangler without any chords passing through origin transforms a pair of orthonormal product states into a pair of orthonormal maximally entangled states?
@@
In this paper, we provide answer to this question. First, we identify the regions of perfect entanglers with at least one chord passing through origin. Next, we derive two sets of orthogonal conditions between a pair of product states obtained from the argand diagram of perfect entanglers without any chord passing through origin. We express them in terms of eigenvalues of generators of perfect entanglers. Satisfying one of these conditions is both necessary and sufficient for a perfect entangler without any chord passing through origin to transform a pair of orthonormal product states into maximally entangled states. Finally, we show that in addition to entangling power, the chords present in the argand diagram can also be used to quantify the gate typicality and linear entropy [6] of two-qubit gates.
@@ II Background
A two-qubit gate $U\in\textnormal{SU}(4)$ can be decomposed as follows [5].
@@
where $L_{1},~{}L_{2}\in\textnormal{SU}(2)\otimes\textnormal{SU}(2)$ are local parts of $U$ and $U_{d}(c_{x},c_{y},c_{z})$ is the nonlocal part of $U$. Two-qubit gates having the same nonlocal part but differing in the local parts form a local equivalence class and each local equivalence class of two-qubit gates is geometrically represented as a point of tetrahedron shown in Fig. 1.
@@
Figure 1: Geometry of local equivalence classes of two-qubit gates. Region coloured in blue is the region of perfect entanglers. The line in red colour represent special perfect entanglers.
@@
$U_{d}(c_{1},c_{2},c_{3})=\sum_{j=1}^{4}e^{ih_{j}/2}|\Psi_{j}\rangle\langle\Psi_{j}|,$ (5)
where the column vectors $|\Psi_{j}\rangle$ form the eigen basis of $U_{d}$. It is also called as magic basis [5] and their expressions are given below.
$|\Psi_{1}\rangle=\frac{1}{\sqrt{2}}\left[\begin{tabular}[]{c}1\\ 0\\ 0\\ 1\end{tabular}\right],~{}|\Psi_{2}\rangle=\frac{1}{\sqrt{2}}\left[\begin{tabular}[]{c}0\\ i\\ i\\ 0\end{tabular}\right],$ (14)
$|\Psi_{3}\rangle=\frac{1}{\sqrt{2}}\left[\begin{tabular}[]{c}0\\ 1\\ -1\\ 0\end{tabular}\right],~{}|\Psi_{4}\rangle=\frac{1}{\sqrt{2}}\left[\begin{tabular}[]{c}i\\ 0\\ 0\\ -i\end{tabular}\right].$ (23)
@@
$c_{x}+c_{y}\geq\frac{\pi}{2},$
$c_{y}\pm c_{z}\leq\frac{\pi}{2}.$ (24)
The region of perfect entanglers (coloured in blue) is shown in Fig. 1. Special perfect entanglers which are perfect entanglers with maximum entangling power [9] are represented by the red colour line in Fig. 1. For perfect entanglers, the convex hull of squared eigenvalues of $U_{d}$ encloses zero. The argand diagram of a typical perfect entangler (non perfect entangler) is shown in Fig. 2a (Fig. 2b).
@@
III Perfect entanglers with at least a chord passing through origin
@@
In this section, we identify the regions of perfect entanglers with at least one chord passing through origin. The chord connecting $e^{ih_{j}}$ and $e^{ih_{k}}$ passes through origin only if $|h_{j}-h_{k}|=\pi$. It can be verified that the equations $|h_{j}-h_{j+1}|=\pi$ with $j=1,2,3$ describe the three boundary planes separating perfect entanglers from non perfect entanglers inside the Weyl chamber [Eq. 24]. These three planes are shown in Fig. 3a. Similarly, the equations $|h_{j}-h_{j+2}|=\pi$ with $j=1,2$ describe the planes $c_{x}\pm c_{z}=\pi/2$ [Fig. 3b]. These two planes together form one of the reflecting plane, $c_{x}+\textnormal{sign}(c_{z})c_{z}=\pi/2$, describing the mirror operation [17]. The equation $|h_{1}-h_{4}|=\pi$ which can be rewritten as $c_{x}-c_{y}=\pi/2$ is satisfied only by the point representing CNOT equivalence class. Thus only the gates represented by the five planes shown in Fig. 3a and Fig. 3b have at least one chord passing through origin and hence they can transform a pair of orthonormal product states into a pair of orthonormal maximally entangled states.
@@
All the gates represented by a specific plane have unique argand diagram with a specific chord passing through origin. However, there are gates present in more than one plane. SPEs are at the intersection of $|h_{j}-h_{j+2}|=\pi~{}(j=1,2)$ planes. Hence, for SPEs, two chords connecting diametrically opposite points $(e^{ih_{j}}~{}\textnormal{and}~{}e^{ih_{j+2}}~{}\textnormal{with}~{}j=1,2)$ pass through origin and they can transform orthonormal product basis into orthonormal maximally entangled basis [13, 9]. Among SPEs, the point representing iSWAP equivalence class is also present in $|h_{j}-h_{j+1}|=\pi~{}\textnormal{with}~{}j=1,3$ planes. Similarly, the point representing CNOT equivalence class also satify the conditions, $|h_{2}-h_{3}|=\pi$ and $|h_{1}-h_{4}|=\pi$. Hence, these two equivalence classes have argand diagrams with four chords passing through origin [13].
@@
Similar to SPEs, there are local equivalence classes at the intersection of two planes: between the pairs of planes $\{|h_{3}-h_{4}|=\pi,|h_{1}-h_{3}|=\pi\}$, $\{|h_{2}-h_{3}|=\pi,|h_{1}-h_{3}|=\pi\}$, $\{|h_{1}-h_{2}|=\pi,|h_{2}-h_{4}|=\pi\}$, and $\{|h_{2}-h_{3}|=\pi,|h_{2}-h_{4}|=\pi\}$. However, for these equivalence classes, the two chords connecting diametrically opposite points do not pass through origin but, only one of them passes through origin. Hence these equivalence classes cannot transform orthonormal product basis into orthonormal maximally entangled basis.
@@
The point representing $\sqrt{\textnormal{SWAP}}$ equivalence class is common to $|h_{3}-h_{4}|=\pi$, $|h_{1}-h_{3}|=\pi$, and $|h_{2}-h_{3}|=\pi$ planes. Hence, in the argand diagram of $\sqrt{\textnormal{SWAP}}$ equivalence class, three chords from the point $e^{ih_{3}}$ pass through origin. It implies that the remaining three points $e^{ih_{1}}$, $e^{ih_{2}}$, and $e^{ih_{4}}$ coincide with each other and the argand diagram has only three chords. Similarly, it can be verified that the argand diagram of $\sqrt{\textnormal{SWAP}}^{\dagger}$ equivalence class has only three chords from the point $e^{ih_{2}}$ passing through origin. Thus for the gates belonging to these two equivalence classes there exist three different pairs of orthonormal product states that can be transformed into pairs of orthonormal maximally entangled states.
@@
Figure 3: Regions of perfect entanglers with at least one chord passing through origin. In the subfigure (a) the planes coloured in magenta, green, and yellow are $|h_{1}-h_{2}|=\pi$, $|h_{2}-h_{3}|=\pi$, and $|h_{3}-h_{4}|=\pi$ planes respectively. The violet, purple, and orange lines (excluding the end points) represent perfect entanglers without any chord passing through origin but capable of transforming a pair of orthonormal product states into a maximally entangled states. These perfect entanglers satisfy the conditions obtained in case 2. In the subfigure (b), the planes coloured in red and blue are $|h_{1}-h_{3}|=\pi$ and $|h_{2}-h_{4}|=\pi$ planes respectively.
IV Perfect entanglers without any chord passing through origin
@@
In this section, we consider the perfect entanglers without any chord passing through origin. The argand diagram of such a perfect entangler is shown in Fig. 2a. The origin can be expressed as convex combination of at most three points [18]. In Fig. 2a, origin is contained in two triangles: the triangle with vertices $e^{ih_{j}}$, $e^{ih_{l}}$, and $e^{ih_{m}}$, and the triangle with vertices $e^{ih_{k}}$, $e^{ih_{l}}$, and $e^{ih_{m}}$. Hence, there exist two sets of weights or co-ordinates such that
@@
with $|\phi_{j}|^{2}+|\phi_{l}|^{2}+|\phi_{m}|^{2}=1$ and $|\phi^{\prime}_{k}|^{2}+|\phi^{\prime}_{l}|^{2}+|\phi^{\prime}_{m}|^{2}=1$. From Eqs. 25 and 26, following two product states can be constructed [5]
$|\Phi_{1}\rangle=|\phi_{j}|e^{-i(a\pi+h_{j}/2)}|\Psi_{j}\rangle+|\phi_{l}|e^{-i(b\pi+h_{l}/2)}|\Psi_{l}\rangle$
$+|\phi_{m}|e^{-i(c\pi+h_{m}/2)}|\Psi_{m}\rangle,$ (27)
@@
$|\Phi_{2}\rangle=|\phi^{\prime}_{k}|e^{-i(d\pi+h_{k}/2}|\Psi_{k}\rangle+|\phi^{\prime}_{l}|e^{-i(f\pi+h_{l}/2)}|\Psi_{l}\rangle$
$+|\phi^{\prime}_{m}|e^{-i(g\pi+h_{m}/2)}|\Psi_{m}\rangle,$ (28)
for some integers $a,b,c,d,f,$ and $g$. Now we consider the following cases.
Case 1: First, we consider the condition for the two product states belonging to two different triangles (Eqs. 27 and 28) to be orthogonal. This condition can be written as follows.
Choosing $(f-b)$ as even (odd) and $(g-c)$ as odd (even), above equation can be rearranged as follows.
$\frac{|\phi_{l}|}{|\phi_{m}|}=\frac{|\phi^{\prime}_{m}|}{|\phi^{\prime}_{l}|}.$ (30)
This is satisfied when the following condition is satisfied.
$\frac{|\phi_{l}|^{2}}{|\phi_{m}|^{2}}=\frac{|\phi^{\prime}_{m}|^{2}}{|\phi^{\prime}_{l}|^{2}}.$ (31)
The weight $|\phi_{l}|^{2}$ can be written as the ratio of the area of subtriangle with vertices $[\cos(h_{j}),\sin(h_{j})]$, $[0,0]$, and $[\cos(h_{m}),\sin(h_{m})]$ to the area of triangle with vertices $[\cos(h_{j}),\sin(h_{j})]$, $[\cos(h_{l}),\sin(h_{l})]$, and $[\cos(h_{m}),\sin(h_{m})]$ [19]. Similarly, other weights can also be expressed and the condition for the product states given in Eqs. 27 and 28 to be orthogonal can be expressed in terms of the eigenvalues of $H$ as follows.
$\frac{|\sin(h_{j}-h_{m})|}{|\sin(h_{l}-h_{j})|}=\frac{|\sin(h_{l}-h_{k})|}{|\sin(h_{k}-h_{m})|}.$ (32)
Case 2: Now, we consider the case where $e^{ih_{j}}$ coincides with $e^{ih_{k}}$ in Fig. 2a. In this case, we have $|\phi^{\prime}_{k}|=|\phi_{j}|$, $|\phi^{\prime}_{l}|=|\phi_{l}|$, and $|\phi^{\prime}_{m}|=|\phi_{m}|$ in Eqs. 26 and 28. Choosing $(f-b)$ as even (odd) and $(g-c)$ as odd (even), the condition for the two product states (Eqs. 27 and 28) to be orthogonal becomes
$|\phi_{l}|^{2}=|\phi_{m}|^{2}.$ (33)
$\frac{|\sin(h_{j}-h_{m})|}{|\sin(h_{l}-h_{j})|}=1.$ (34)
Case 3: Now, we derive the condition for the existence of a pair of orthonormal product states with respect to a triangle enclosing the origin. For example, in Fig. 2a, we consider the triangle with vertices $e^{ih_{j}}$, $e^{ih_{l}}$, and $e^{ih_{m}}$. A product state corresponding to this triangle is given in Eq. 27. We construct another product state with respect to the same triangle, for some integers $a^{\prime},b^{\prime},~{}\textnormal{and}~{}c^{\prime}$, as shown below.
$|\Phi^{\prime}_{1}\rangle=|\phi_{j}|e^{-i(a^{\prime}\pi+h_{j}/2)}|\Psi_{j}\rangle+|\phi_{l}|e^{-i(b^{\prime}\pi+h_{l}/2)}|\Psi_{l}\rangle$
$+|\phi_{m}|e^{-i(c^{\prime}\pi+h_{m}/2)}|\Psi_{m}\rangle.$ (35)
@@
Along with the condition $|\phi_{j}|^{2}+|\phi_{l}|^{2}+|\phi_{m}|^{2}=1$, it can be found that $|\phi_{j}|^{2}=1/2$. This condition can be written in terms of eigenvalues of $H$ as follows.
$\frac{2|\sin(h_{m}-h_{l})|}{|\sin(h_{l}-h_{j})+\sin(h_{j}-h_{m})+\sin(h_{m}-h_{l})|}=1.$ (38)
This condition will be satisfied only if the following equation is satisfied.
$\sin(h_{l}-h_{j})+\sin(h_{j}-h_{m})=\sin(h_{m}-h_{l})$ (39)
Taking $h_{l}-h_{j}=\theta_{1}$ and $h_{j}-h_{m}=\theta_{2}$ for some $\theta_{1}$ and $\theta_{2}$ which could be either positive or negative, we get
$\sin(\theta_{1})+\sin(\theta_{2})=-\sin(\theta_{1}+\theta_{2})$ (40)
This condition is satisfied when either $\theta_{1}=\pm\pi$ or $\theta_{2}=\pm\pi$. This is also satisfied when $\theta_{1}=\pm\pi$ and $\theta_{2}=\pm\pi$, but in this case, $e^{ih_{l}}$ and $e^{ih_{m}}$ coincide with each other. Thus, the weight $|\phi_{j}|^{2}$ to be 0.5, one of the chords from $e^{ih_{j}}$ should pass through origin. Thus, to construct a pair of orthogonal product states (which can be converted into maximally entangled states) with respect to a triangle containing origin, the origin should be at one of the edges of the triangle. So, this case covers only the perfect entanglers with a chord passing through origin.
Now, we consider two product states, $|\xi\rangle=\sum_{j}\epsilon_{j}|\Psi_{j}\rangle$ and $|\xi^{\prime}\rangle=\sum_{j}\epsilon^{\prime}_{j}|\Psi_{j}\rangle$ $(j=1,2,3,4)$ in the magic basis with the conditions $\sum_{j}|\epsilon_{j}|^{2}=\sum_{j}|\epsilon^{\prime}_{j}|^{2}=1$ and $\sum_{j}\epsilon_{j}^{2}=\sum_{j}\epsilon^{\prime 2}_{j}=0$ [5]. If these two product states are transformed into maximally entangled state by $U_{d}$ [Eq. 5] with $h_{j}-h_{k}\neq\pi$ for all $j\neq k$, then we should have (up to a global phase) $\epsilon_{j}=\pm|\epsilon_{j}|e^{-ih_{j}/2}$ and $\epsilon^{\prime}_{j}=\pm|\epsilon^{\prime}_{j}|e^{-ih_{j}/2}$ for all $j$ [5]. In addition, we should have $\sum_{j}|\epsilon_{j}|^{2}e^{-ih_{j}}=\sum_{j}|\epsilon^{\prime}_{j}|^{2}e^{-ih_{j}}=0$ [5]. Since, the squared eigenvalues of $U_{d}$ ($e^{-ih_{j}}$) exist on a plane, one of the weights should be zero in both summations [18]. It can be verified that if the product states, $|\xi\rangle$ and $|\xi^{\prime}\rangle$, are orthogonal then the argand diagram of $U_{d}$ belongs to the kinds which satisfy one of the conditions obtained in cases 1 and 2.